# UINav: A Practical Approach to Train On-Device Automation Agents

### v2

@@ Abstract
Wei Li, Fu-Lin Hsu, Will Bishop, Folawiyo Campbell-Ajala, Max Lin, Oriana Riva Google Research
**footnotetext: Worked as an intern at Google Research
@@ 6.1 Agent and referee accuracy
Table 2 reports the results. For ablation purposes, we consider three variants of UINav, depending on whether demonstration augmentation (+aug) and utterance masking (+utt) are enabled. UINav+aug surpasses all baselines by 7 and 11 percentage points in task accuracy and 8.5 and 0.7 points in step accuracy. Without demo augmentation UINav outperforms all three baselines, in all except one case (step accuracy in app unseen and task seen). This demonstrates the effectiveness of the UINav design and how demo augmentation effectively exposes the model to a larger variety of training conditions thus improving generalizability. Generalizing to new apps is generally harder than generalizing to new tasks, but with the addition of utterance matching UINav achieves 59.6% in task accuracy and 81.9% in step accuracy, well above all baselines.
@@ 6.2 Demonstration effort
We collect demonstrations with the goal to achieve near perfect success rates. With the exception of the search task we collect from 10 to 106 demonstrations (on average 32.7) per task, 3661 in total (Fig. 4). Collecting 10 demonstrations takes less than 10 minutes. The search task must work across 100+ apps hence requiring 1700+ samples. To verify this data is sufficient to train accurate agents, in a second phase we collect additional 596 test samples. Because of the random initialization of the environment, and the dynamic characteristics of a live system, it is unlikely that the models see a training sample that is identical to a test one. The UINav agent achieves 90.6% task accuracy and 95.8% step accuracy; the referee is 99.5% accurate.
@@ 7 Conclusions
Appendix A Appendix
Limitations
To limit the number of required demonstrations, the UINav agent makes decisions based only on the contents of the current screen and does not utilize information from previous screens. However, if a task truly requires an agent to remember previous states or actions, then the current architecture of the agent model will fail. Our assumption is that a well-designed user interface often presents all the information that is needed for human action on the current screen. Fortunately, the accuracy of our memory-less agent proves that this assumption is true for all the tasks we have tested so far. If in the future, UINav runs into tasks that require memory, the agent model can be enhanced with memory through either a recurrent neural network or padding previous states in its input.
@@ A.1 An example task: search in YouTube
Fig. 4 shows the SEARCH task required over 1700 task demonstrations because it must work for 100 or more different apps and websites. All other tasks are specific to a single app and thus require fewer samples, 33 on average.
@@ A.5 Implementation
The neural networks are agnostic to whether the Android Accessibility View Hierarchy or screen understanding techniques are used to produce screen representations. We include demonstrations using both data sources in the same pool of training samples. Both approaches have their limitations. There are icons that are unrecognizable by the icon detectors of screen understanding models and the output of text recognizer may contain errors. On the other hand, visible UI elements may be absent in the corresponding Accessibility tree if the app contains Web views, Canvas, etc.

### v3

@@ Abstract
Wei Li22footnotemark: 2   Fu-Lin Hsu33footnotemark: 3thanks: Work done as an intern at Google Research.   Will Bishop22footnotemark: 2  Folawiyo Campbell-Ajala22footnotemark: 2   Max Lin22footnotemark: 2   Oriana Riva22footnotemark: 2 22footnotemark: 2  Google Research 33footnotemark: 3  University of Pennsylvania
@@ 6.1 Agent and referee accuracy
Table 2 reports the results. For ablation purposes, we consider three variants of UINav, depending on whether demonstration augmentation (+aug) and utterance masking (+utt) are enabled. UINav+aug surpasses all baselines by 7 and 11 percentage points in task accuracy and 8.5 and 0.7 points in step accuracy. Without demo augmentation UINav outperforms all three baselines, in all except one case (step accuracy in app unseen and task seen). This demonstrates the effectiveness of the UINav design and how demo augmentation effectively exposes the model to a larger variety of training conditions thus improving generalizability. In this dataset, generalizing to new apps appears to be harder than generalizing to new tasks. With the addition of utterance matching, on unseen apps, UINav still achieves 59.6% in task accuracy and 81.9% in step accuracy, well above all baselines.
@@ 6.2 Demonstration effort
We collect demonstrations with the goal to achieve near perfect success rates. With the exception of the search task we collect from 10 to 106 demonstrations (on average 32.7) per task, 3661 in total (Fig. 4). Collecting 10 demonstrations takes less than 10 minutes. The search task must work across 100+ apps hence requiring 1700+ samples. To verify this data is sufficient to train accurate agents, in a second phase we collect additional 596 test samples. Because of the random initialization of the environment, and the dynamic characteristics of a live system, it is unlikely that the models see a training sample that is identical to a test one. The UINav agent achieves 90.6% task accuracy and 95.8% step accuracy; the referee is 99.5% accurate. Please note that the numbers of demonstrations in Fig. 4 are most likely more than the minimum required to reach the same accuracy, as we prioritize improving accuracy over reducing the number of training samples. It is less effort to add new demonstrations as a batch than finding out whether a specific demonstration improves model accuracy.
In an informal user study, a few software engineers with no prior experience using UINav utilized it to build agents for a few tasks. They started from scratch, without using any existing demonstrations. The time spent on collecting data for each task was between 10 to 20 minutes while all participants claimed their resulting agents performed perfectly.
@@
7 Limitations
To limit the number of required demonstrations, the UINav agent makes decisions based only on the contents of the current screen and does not utilize information from previous screens. However, if a task truly requires an agent to remember previous states or actions, then the current architecture of the agent model will fail. Our assumption is that a well-designed UI often presents all the information that is needed for successful human interaction on the current screen. The accuracy of our memory-less agents proves that this is the case for the tasks tested so far. For tasks or UIs that require memory, the UINav agent model can be enhanced with memory through either a recurrent neural network or by padding previous states in its input.
@@ A.1 An example task: search in YouTube
Fig. 4 shows the SEARCH task requires over 1700 task demonstrations because it must work for 100 or more different apps and websites. All other tasks are specific to a single app and thus require fewer samples, 33 on average.
@@ A.5 Implementation
The selection of SmallBERT over a larger language model is mainly for on-device inference. We restrict the input utterances to predefined patterns so that arguments can be parsed through regular expressions. With the help of utterance masking, our models deal with higher data diversity and maintain high-accuracy. If an LLM can be used, such restrictions won’t be necessary.
@@
The neural networks are agnostic to whether the Android accessibility tree or screen understanding techniques are used to produce screen representations. We include demonstrations using both data sources in the same pool of training samples. Both approaches have their limitations. There are icons that are unrecognizable by the icon detectors of screen understanding models and the output of text recognizer may contain errors. On the other hand, visible UI elements may be absent in the corresponding accessibility tree if the app contains Web views, Canvas, etc.